OpenStudy (anonymous):

What value does P need to have in order for x^2 4x + P to be a perfect square?

OpenStudy (helder_edwin):

remember this \[ \large (a\pm b)^2=a^2\pm2ab+b^2 \]

OpenStudy (anonymous):

ya

OpenStudy (lgbasallote):

also remember this \[\left(\frac b2\right)^2 = c\] when you have a quadratic equation like \[ax^2 + bx + c\]

OpenStudy (anonymous):

no on igbasallote

OpenStudy (lgbasallote):

what i mean is \[P = \left(\frac 42 \right)^2\] so just simplify that and you get your answer

OpenStudy (anonymous):

4 thanks